OpenStudy (anonymous):

Which values of x and y satisfy 2x − y = 7 and 3x − 2y = 5 simultaneously? x = 1 and y = -5 x = 1 and y = -1 x = 4 and y = 1 x = 9 and y = 11

OpenStudy (anonymous):

This system of equations appears ideal for solving using the "substitution" method. You can do this by taking the first equation 2x - y = 7 and solve for y, you should get y= 2x-7. Now substitute this for y in the second equation (3x-2y=5) after substitution it would look like this: 3x - 2(2x-7) = 5 which becomes 3x -4x + 14 = 5 which becomes -x =-9 or x = 9. Now plug that value for x into the second equation and solve for y.

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

thank you